Assume that you would like to purchase 100 shares of preferred stock that pays an annual dividend of $6.00 per share. However, you have limited resources now, so you cannot afford the purchase price. In fact, the best you can do now is to invest your money in a mutual fund that offers an average return of 6% compounded monthly. Because the preferred stock is riskier, it has an annual rate of return of 12% (assume that this rate will remain constant into the foreseeable future). For you to be able to purchase this stock at the end of 5 years, how much must deposit in your bank account today?
Answer #1

price of Preferred Stock = 6/0.12

Price of Preferred Stock= $50

Value of purchase = 100(50) = $5,000

Amount in mutual fund = 5,000/(1.06)^5

Amount in mutual fund= $3,736.29